### 12 Jan 2024 v1

# APLe: Token-Wise Adaptive for Multi-Modal Prompt Learning

Abstract
Pre-trained Vision-Language (V-L) models set the benchmark for generalization to downstream tasks among the noteworthy contenders. Many characteristics of the V-L model have been explored in existing research including the challenge of the sensitivity to text input and the tuning process across multi-modal prompts. With the advanced utilization of the V-L model like CLIP, recent approaches deploy learnable prompts instead of hand-craft prompts to boost the generalization performance and address the aforementioned challenges. Inspired by layer-wise training, which is wildly used in image fusion, we note that using a sequential training process to adapt different modalities branches of CLIP efficiently facilitates the improvement of generalization. In the context of addressing the multi-modal prompting challenge, we propose Token-wise Adaptive for Multi-modal Prompt Learning (APLe) for tuning both modalities prompts, vision and language, as tokens in a sequential manner. APLe addresses the challenges in V-L models to promote prompt learning across both modalities, which indicates a competitive generalization performance in line with the state-of-the-art. Preeminently, APLe shows robustness and favourable performance in prompt-length experiments with an absolute advantage in adopting the V-L models.
Figure 1: Comparison of framework. CoOp adopts a uni-modal prompting. MaPLe demonstrates prompt learning in a multi-modal manner by the coupling function. APLe proposes an independent and sequential multi-modal prompt learning with adaptation.
1 Introduction
Vision-language (V-L) pretraining is an emerging paradigm that performs image recognition tasks, such as image-recognition recognition1 (9, 19, 41), object detection object-detection1 (8, 26, 40, 35), and segmentation segmentation1 (6, 22, 31), inspirational elevating the performance in generalization ability. A novel model, CLIP CLIP (30) indicates the success of the paradigm by leveraging contrastive learning to align large-scale text-image pairs. This model introduces a simple inference process that pairs text embedding from a hand-crafted query and image embedding by calculating cosine similarity to predict the class of the image, which demonstrates the excelling generalization capability.
CLIP triggers the generalization toward a new era with potential. An earlier study, CoOp COOP (44), introduced a learnable continuous prompt approach as a text input instead of relying on hand-crafted prompts shown in Fig. 1. CoOp further reveals the potential and challenge of the vision-language pretraining paradigm by delivering exceptional results on the accuracy of image classification. It also tackles the challenges of the extensive time demanded for training. Recent-coming research, MaPLe multi-modal (17), addresses the insufficient generalization performance for the downstream task if only the language prompt is explored. Thus, MaPLe exploits the prompting approach in both modalities’ prompts demonstrated in Fig. 1. The prompts are learned from both vision and language sources, while MaPLe initiates vision prompts from language counterparts by the unidirectional parsing function.
Despite the effectiveness of MaPLe toward sufficiently exploiting both behaviours of both modalities to contrastive learning. Its parsing mechanism makes it infeasible to conduct prompt learning in two modalities equally due to the disparate sophistication and conditions. Such mechanisms potentially introduce challenging problems resulting in generalization capability, such as overfitting caused by the prompt length. To this end, inspired by layer-wise training and image recognition that contributes to separately extracting and processing the features from different images. This paper proposes Token-wise Adaptive for Multi-modal Prompt Learning (APLe) to adopt multi-modal prompt learning independently and sequentially to tackle the challenges and boost generation capability with the framework shown in Fig. 1.
In the fundamental generalization evaluation experiments, APLe demonstrated a comparable performance with the existing methods, particularly in datasets with domain shifts. Distinctively, APLe exploits the knowledge from both modalities and illustrates the advantage of robustness by sequential prompt learning and image adaptation, while the existing approaches are insufficient or omit the image conditions and multi-modal prompt conditions. Thanks to its effectiveness, APLe attains competitive generalization capability over the baselines owing to the effective addressing of the challenges caused by image complexity and prompt length. To sum up, the contributions of the paper are summarized as follows:
  • We propose the token-wise adaptive for multi-modal prompt learning to adopt the V-L model. To the best of our knowledge, this is the first approach to address the overfitting caused by the prompt length and the disparate sophistication of features caused by multiple modalities.
  • To extend the generalization capability, The paper proposes the token-wise knowledge training framework that allows a prompt learning process to be incorporated with CLIP zero-shot knowledge in a prompt-wise manner. Such a framework promotes APLe to learn the knowledge from two modalities without interference, which attains adaptive and stable generalization capability.
  • This approach conducts prompt learning with zero-shot knowledge followed by a token adaptation function to alleviate knowledge conflicts and encourage collaboration between modalities. The adaptation function encourages APLe to align the V-L representation.
Figure 2: Overview of APLe (Token-wise Adaptive for Multi-modal Prompt Learning) framework for prompting. The prompts have been trained sequentially as tokens with CLIP zero-shot knowledge and adaptation to alleviate the knowledge conflicts and prompt the synergy between modalities. APLe also deploys an image adapter to mitigate image noise and enhance image features effectively. Specifically, CE denotes the cross-entropy loss, KL denotes the Kullback-Leibler loss and $\lambda$ presents the hyper-parameter used to combine the loss functions.
2 Related Work
Vision Language Models: The multi-modalities research in exploring vision-language interaction is a core topic. At the early stage, a few attention-based approaches are introduced including BAN BAN (18), Intra-Inter Intra-Inter (10) and MCAN MCAN (39), which make contributions to the downstream tasks. While the following approaches are proposed with BERT-like architecture, ViLBERT ViLBERT (25), LXMERT LXMERT (34) and UNITER UNITER (3), expend the boundary in reasoning between the modalities. In the near past, the two prominent models, CLIP CLIP (30) and ALIGN ALIGN (14) have dominated the visual-language tasks through disparate representation learning with an impressive performance.
To adopt the V-L model in downstream tasks, the conventional approaches do not adapt the current preference, which constrains the zero-shot or few-shot capability by fine-tuning and linear probing. Thanks to the supervised training with abundant data from the internet, $\sim$400M image-text pairs used in CLIP and $\sim$1B used in ALIGN, both models demonstrated the power of visual-language pre-training paradigm in zero-shot learning and few-shot learning. Inspired by CLIP and ALIGN, many following approaches achieve dominant generalization performance on specific tasks by adopting the V-L paradigm including few-shot image-recognition recognition1 (9, 19, 41), object detection object-detection1 (8, 26, 40) and segmentation segmentation1 (6, 22, 31).
Prompt Design: The prompt design inspires research interest by the brand new paradigm, “pre-train, prompt, and predict”, which was used in NLP tasks NLP1 (21, 23, 24) at the early stage. The recently proposed approaches introduced the prompt learning paradigm to adopt the V-L model to downstream tasks which takes a query, as known as prompt, into the language branch to boost the generalization capability in understanding tasks. Therefore, by the effectiveness of prompt design, a rising trend of research lies on prompt design or prompt learning to further extend the potential of the V-L models and address the newly arisen challenges caused by prompt learning such as the language-only prompting COOP (44, 43, 45), vision-only prompting vision1 (15, 36, 37) and a recently proposed method, MaPLe, in multi-modal prompting multi-modal (17).
Prompt Learning in Vision and Language: The early work CoOp COOP (44) exploits a learnable token into a language branch for fine-tuning CLIP. Its subsequent work Co-CoOpCOCOOP (43) further extends the boundary of generalization by exploiting conditional prompt. For the works applying prompt learning in the vision branch, Visual1 (1) is the first work to perform vision prompt in CLIP while video18 (16) introduces the prompt learning paradigm into video understanding tasks. The recent work, MaPLe multi-modal (17), proposed a framework that applies learnable tokens into both vision and language branches by coupling function. MaPLe shows an advanced performance in downstream tasks owing to the coupling function improving the alignment between vision and language representations.
3 Method
The proposed approach tackles the V-L pretraining model, CLIP, in downstream tasks. The holistic framework of our approach, APLe, is demonstrated in Fig 2. We argue that the two modalities, vision and language, equally contribute to downstream tasks while also delving into the challenges of disparate sophistication in prompt learning. To efficiently address the challenges,APLe proposes a framework that encourages prompts learned independently and sequentially from two modalities. Explicitly, APLe first attaches a Gaussian Filter image adapter to alleviate the sophistication of the image features before prompt learning, which empowers the model to effectively mitigate image noise and enhance image features across various scales. Then, APLe is structured into two blocks. A sequential token-wise knowledge training block first explores learnable prompts as tokens for training with the CLIP zero-shot model, which obtains generalization knowledge and encourages multi-modality tokens to learn knowledge independently and sequentially without interference from the other. Ultimately, the multi-modal token adaptation block promotes the synergy between the tokens.
3.1 Revisiting CLIP
CLIP presents a basic architecture to adopt the vision-language (V-L) model and addresses image recognition and classification tasks. Many existing approaches follow the architecture of CLIP COOP (44, 41). Many components used in CLIP including the neural network backbone, transformer and Vit (vision transformer), are inherited in the recent research and our approach, APLe. To adapt feature representation in CLIP, the text encoder transformed a hand-crafted query $Q$ as prompt into feature representations, and the image encoder progressed likewise to the image as features $I\in\mathbb{R}^{3\times H\times W}$ for cosine similarity calculation and generating the prediction result. The process pipeline for CLIP is demonstrated below.
Text Feature Block: Initially, the text query $Q$ is tokenized and transits into word embedding. Then, the transformer layers $\mathcal{B}$ compute word embedding and forward its result to the projection function for the cosine similarity calculation.
$W=Embed(Tokenizer(Q))\qquad W\in\mathbb{R}^{T\times d_{lang}}$ (1)
$W_{n}=\mathcal{B}_{n}(W_{n-1})\qquad n=1,2,3...,K.$ (2)
$Z_{K}=TextProj(W_{K})\qquad Z\in\mathbb{R}^{d_{vis-lang}}$ (3)
Image Feature Block: Similar to text feature processing, the procedure excludes tokenization but appends a concatenation function for combining the class identification token $c$.
$E=PatchEmbed(I)\qquad$ (4)
where E $\in\mathbb{R}^{M^{2}\times{d_{vis}}}$ with the fixed-size patches M = 14 .
$\left[c_{n},E_{n}\right]=\mathcal{V}_{n}([c_{n-1},E_{n-1}])\qquad n=1,2,3...,K.$ (5)
where $c$ denote as the token for class identification, [,][\;,\;][ , ] as concatenation function, $\mathcal{V}$ as vision transformer layers.
$f=ImageProj(c_{k})\qquad f\in\mathbb{R}^{d_{vis-lang}}$ (6)
Where $c_{k}$ is produced in $K$ steps in Equation. 5. The result is transferred into the common V-L latent space by $ImageProj$ projection function for the final operation in Equation. 7.
Overall: The text query in the template “a photo of {class}” is transferred as $Z_{i}$, the feature $f$ of the target image and the class to predict denoted as $i\in$1,2,3…C}, the final step for cosine similarity calculation and predictions shown below:
$p_{i}=\frac{\mathrm{exp}(sim(Z_{i}\cdot f)/\tau)}{{\textstyle\sum_{j=1}^{C}}\mathrm{exp}(sim(Z_{j}\cdot f))}$ (7)
where $\tau$ stands for the temperature parameter, $sim(\cdot)$ denoted as cosine similarity calculation and $p_{i}$ as the prediction result.
3.2 APLe: Token-wise knowledge training and Multi-modal Token Adaptation
This paper argues that the existing methods for uni- or multi-modalities carry out the learnable prompt training with weak considering the potential and challenge of the differences in vision and language modalities. To this end, we introduce token-wise training and adaptation framework to multi-modalities prompt learning which addresses the challenges by independently and sequentially training for both prompts and optimizing the utilization of the potential of the V-L model. Image Fusion with Adapter: We argue that disparate sophistication in images caused difficulty in prompt learning. To this end, a simple and commonly seen function to obtain a lucid and authentic image feature from the original image. Inspired by the existing work in image recognition tasks (e.g. CDDFuse CDD (42)), the image $I$ in APLe is processed by Fast Fourier Transform $\mathcal{F}(\cdot)$ followed by the Gaussian filter $\mathcal{G}(\cdot)$ for suppressing image noise and enhancing subtle image features. the simple processing pipeline is shown below.
$\mathcal{G}(x,y)$ $=\frac{1}{2\pi\sigma^{2}}\cdot e^{-\frac{x^{2}+y^{2}}{2\sigma^{2}}}$ (8)
$I^{{}^{\prime}}$ $=\mathcal{G}(\mathcal{F}(I))$ (9)
Here $\sigma$ refers to the standard deviation in $\mathcal{G}(\cdot)$ and is set to 0.05 in our case, $x$ to the x-coordinate of each pixel in the image and $y$ to the y-coordinate of each pixel in the image.
$\bar{I}=\alpha I+(1-\alpha)\cdot\mathcal{F}^{-1}(I^{{}^{\prime}})$ (10)
Then, the image fusion operation is applied to the adjusted image and the original image where $\alpha$ is set to 0.9 The result in image fusion $\bar{I}$ is sequentially taken by Eqs.15-17 to produce the image final feature for contrastive learning. The fundamental aim of image fusion is to stabilize the variance of image features by suppressing interference rather than eliminating it. While image fusion alleviates the complexity of the image feature, the challenge is further efficiently compensated by vision tokens training in the APLe token-wise knowledge training and adaptation framework, which allows APLe to enhance its generalization performance further.
$p_{i}=\frac{\mathrm{exp}(sim(Z_{i}\cdot\bar{f})/\tau)}{{\textstyle\sum_{j=1}^{C}}\mathrm{exp}(sim(Z_{j}\cdot\bar{f}))}$ (11)
By introducing fusion image $\bar{I}$ into the image feature embedding pipeline, the new image feature $\bar{f}$ is taken into Eq. 11 for contrastive learning with text feature $Z$ for predicting $p_{i}$.
Feature Embedding: In contrast to the existing approach, APLe initiates learnable prompt $D$ and $G$ for text and image respectively in a decoupling manner, which is embedded into their modalities feature by concatenation operation as seen in the Eq. 13 and Eq. 16. As the effectiveness of deep prompting, the component is maintained to APLe, which is deployed to process the newly concatenated prompts. In other settings, the length $m$ for both learnable tokens is set to 2 and the number of deep transformer layers is set with extra depth $S$. The feature embedding process is shown in the following section.
Feature Embedding in Text Feature: The encoding to text is similar to CLIP, but a learnable token $D$ is embedded.
$W=Embed(Tokenizer(Q))\qquad W\in\mathbb{R}^{T\times d_{lang}}$ (12)
$\begin{split}&[W_{s},D_{s}]=\mathcal{B}_{j}([W_{s-1},D_{s-1}])\\ &s=S+1,...,K.\qquad D\in\mathbb{R}^{m\times d_{lang}}\end{split}$ (13)
$Z_{K}=TextProj(W_{K})\qquad Z\in\mathbb{R}^{d_{vis-lang}}$ (14)
Feature Embedding in Image Feature: The image encoding corresponds to CLIP, but the image is first processed by the image adapter followed by the embedding process. Then, the result is concatenated with the learnable token $G$.
$E=PatchEmbed(I)$ (15)
$\begin{split}&[c_{s},E_{s},G_{s}]=\mathcal{V}_{j}([c_{s-1},E_{s-1},G_{s-1}])\\ &s=S+1,..,K.\qquad G\in\mathbb{R}^{m\times d_{vis}}\end{split}$ (16)
$f=ImageProj(c_{k})\qquad f\in\mathbb{R}^{d_{vis-lang}}$ (17)
In contrast with other approaches, the decoupling structure enables both vision and language prompts to learn knowledge separately and alleviate the constraint from the other, which enhances the capability of multi-modal prompts learning.
Training and Adaptation: Preeminently, APLe explicitly and effectively boost the generalization capability by conducting a distinctive and sequential training and adaptation process for prompt learning across the multiple modalities.
Sequential Token-Wise Knowledge Training: APLe train each modality sequentially and separately as embedding tokens at a time while keeping another one frozen. During the training, APLe incorporates CLIP zero-shot knowledge in the prompt learning to mitigate the potential of overfitting caused by prompt setting and escalates generalization capability using diverse combination parameters to vision and language modalities, $\lambda_{g}$ and $\lambda_{d}$ respectively, as shown in Eqs. 1819.
$Stage1:\quad\mathcal{L}^{I}_{total}$ $=\mathcal{L}_{ce}(p_{i},C_{i})+\lambda_{d}\mathcal{L}_{kl}(p_{i},p_{i}^{zs})$ (18)
$\mathcal{L}^{I}_{total}$ $=\mathcal{L}_{ce}(p_{i},C_{i})+\lambda_{g}\mathcal{L}_{kl}(p_{i},p_{i}^{zs})$ (19)
where $\mathcal{L}_{ce}$ and $\mathcal{L}_{kl}$ are the cross-entropy loss for classification prediction and Kullback-Leibler divergence loss for quantifying the prediction disparity between the trained model and CLIP zero-shot model, which are denoted as $p_{i}$ and $p_{i}^{zs}$ respectively. While separately training the multi-modal prompt, the hyper-parameter $\lambda$ is set for combining with KL loss, while $\lambda_{g}$ and $\lambda_{d}$ are set to 0.3 for the vision prompt training and 0.5 for the language prompt training respectively.
Multi-modal Token Adaptation: Owing to the token-wise knowledge training, APLe explores the potential of both modalities in prompt learning. For further adapting the synergy of the tokens from both branches and attaining advanced generalization capability in downstream tasks, We introduce a straightforward and transparent approach that retrains the model to promote synergy between the tokens without zero-shot knowledge as the end of the pipeline of APLe.
$Stage2:\quad\mathcal{L}^{II}_{total}$ $=\mathcal{L}_{ce}(p_{i},C_{i})\vspace{-0.2cm}$ (20)
4 Experiments
To evaluate the capability of APLe, we conduct experiments with the well-known methods CLIP (30, 44, 43) and one recently proposed method multi-modal (17). The experiment and the datasets used align with prior research including three experiments and datasets of Caltech101 Caltech101 (7), DTD DTD (4), EuroSAT EuroSAT (11), FGVCAircraft FGVCAircraft (27), Flowers102 Flowers102 (28), Food101 Food101 (2), ImageNet ImageNet (5), OxfordPets OxfordPets (29), StanfordCars StanfordCars (20), SUN397 SUN397 (38), UCF101 UCF101 (33), ImageNetSketch object-detection4 (35), ImageNetV2 ImageNetV2 (32), ImageNet-A ImageNet-A (12) and ImageNet-R ImageNet-R (13).
4.1 Fundamental Experiment
Base-to-Novel Generalization: To examine the zero-shot learning capability, the datasets are split into two groups, base classes and novel classes. APLe is trained on the base classes in a few-shot manner and then evaluated in both groups.
Cross-Dataset Generalization: For the evaluation of zero-shot learning and transfer learning, we train APLe in ImageNet datasets by training all 1000 classes in a few-shot manner. Then, the evaluation is conducted in another 10 datasets.
Domain Generalization: To further demonstrate the performance of APLe on generalization capability, The model trained in cross-dataset generalization is evaluated for the other four datasets with a domain shift from ImageNet.
(a) Average over 11 datasets
(b) ImageNet
(c) Caltech101
(d) OxfordPets
(e) StanfordCars
(f) Flowers102
(g) Food101
(h) FGVCAircraft
(i) SUN397
(j) DTD
(k) EuroSAT
(l) UCF101
Name Base Novel HM
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
Co-CoOp 80.47 71.69 75.83
MaPLe 82.28 75.14 78.55
APLe 81.99 75.11 78.40
Name Base Novel HM
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
Co-CoOp 75.98 70.43 73.10
MaPLe 76.66 70.54 73.47
APLe 76.46 70.78 73.51
Name Base Novel HM
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
Co-CoOp 97.96 93.81 95.84
MaPLe 97.74 94.36 96.02
APLe 98.30 94.80 96.52
Name Base Novel HM
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
Co-CoOp 95.20 97.69 96.43
MaPLe 95.43 97.76 96.58
APLe 95.22 97.56 96.37
Name Base Novel HM
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
Co-CoOp 70.49 73.59 72.01
MaPLe 72.94 74.00 73.47
APLe 72.42 74.42 73.41
Name Base Novel HM
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
Co-CoOp 94.87 71.75 81.71
MaPLe 95.92 72.46 82.56
APLe 95.09 73.90 83.17
Name Base Novel HM
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
Co-CoOp 90.70 91.29 90.99
MaPLe 90.71 92.05 91.38
APLe 90.74 91.77 91.25
Name Base Novel HM
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
Co-CoOp 33.41 23.71 27.74
MaPLe 37.44 35.61 36.50
APLe 36.65 34.45 35.52
Name Base Novel HM
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
Co-CoOp 79.74 76.86 78.27
MaPLe 80.82 78.70 79.75
APLe 80.22 78.44 79.32
Name Base Novel HM
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
Co-CoOp 77.01 56.00 64.85
MaPLe 80.36 59.18 68.16
APLe 79.98 57.93 67.19
Name Base Novel HM
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
Co-CoOp 87.49 60.04 71.21
MaPLe 94.07 73.23 82.35
APLe 93.62 74.04 82.69
Name Base Novel HM
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
Co-CoOp 82.33 73.45 77.64
MaPLe 83.00 78.66 80.77
APLe 83.18 78.13 80.58
Table 1: Comparison with existing approaches. APLe demonstrates comparable generalization ability over 11 datasets.
(b) ImageNet
(c) Caltech101
(d) OxfordPets
(e) StanfordCars
(f) Flowers102
(g) Food101
(h) FGVCAircraft
(i) SUN397
(j) DTD
(k) EuroSAT
(l) UCF101
4.2 Base-to-Novel Generalization
Table.1 demonstrates the capability of APLe by comparison among existing prompt learning approaches including original CLIP with zero-shot CLIP (30), CoOp COOP (44), Co-CoOp COCOOP (43) and MaPLe multi-modal (17) over 11 datasets. We evaluate APLe by comparing to Co-CoOp and MaPLe separately to address the diverse characteristics of our approach among the other approaches.
Base Classes in Generalization: While performing prompt learning, Co-CoOp drops the average accuracy of 11 datasets in base classes from 82.69% to 80.47% compared to CoOp. By the independent and sequential framework, APLe stabilized the average accuracy at 81.99% with an advanced generalization capability in all datasets compared to Co-CoOp which demonstrates the capability in line with the baseline.
Novel Classes in Generalization: In novel classes, While APLe applied conditioning effort on the image branch and incorporated zero-shot knowledge, APLe indicated significant improvement over he datasets except for the OxfordPets with the 0.13% drop in novel class. Over 11 datasets, the average accuracy of APLe in novel classes generalization is boosted from 71.69% to 75.11%. Meanwhile, APLe demonstrates a 2.57% significant increase if both base classes and novel classes are considered which is denoted as harmonic means.
DTD FGVCAircraft
Sigma Base Novel HM Base Novel HM
0.02 80.25 60.55 69.02 37.59 33.95 35.68
0.05 79.98 57.93 67.19 36.65 34.45 35.52
0.08 79.59 57.13 66.51 37.13 34.45 35.74
0.10 78.86 57.33 66.39 37.17 35.25 36.19
0.12 79.71 58.77 67.66 37.73 34.93 36.28
MaPLe 80.36 59.18 68.16 37.44 35.61 36.50
Table 2: Effectiveness analysis of sigma in two failure cases.
APLe $vs$ MaPLe: In Table.1 (a), it indicates that the average accuracy is comparable between APLe and MaPLe, MaPLe applied multi-modal prompting with a coupling function. APLe outperforms the advantage in 5/11 datasets and has equivalent performance in the other 4 datasets (average gap = $\sim$0.32%) on average accuracy. The disadvantage occurs at DTD and FGVCAircraft with 1.15% and 1.16% gap respectively. In general, APLe shows competitive performance in 4/11 datasets and comparable results in the other 5 datasets (average gap $\sim$ 0.20%). In DTD and FGVCAircraft, the drop happened with 0.97% and 0.98% respectively.
Source Target
ImageNet Caltech Pets Cars Flowers Food Aircraft SUN DTD EuroSAT UCF Avg.
CoOp 71.51 93.70 89.14 64.51 68.71 85.30 18.47 64.15 41.92 46.39 66.55 63.88
Co-CoOp 71.02 94.43 90.14 65.32 71.88 86.06 22.94 67.36 45.73 45.37 68.21 65.74
MaPLe 70.72 93.53 90.49 65.57 72.23 86.20 24.74 67.01 46.49 48.06 68.69 66.30
APLe 70.55 93.71 90.44 65.98 71.55 86.03 23.54 67.29 46.14 48.85 68.72 66.23
Table 3: Comparison on cross-datasets. APLe shows a competitive result among the methods by the model trained in ImageNet.
Failure Cases: We analyze cases where the APLe fails while MaPLe succeeds. Specifically, Taking into account the non-learnable nature and sensitivity to image style, we assess the effectiveness of different sigma used in the image adapter as shown in Table. 2. The table shows that APLe achieves advanced performance when sigma is set to 0.02 for DTD and 0.12 for FGVCAircraft. Considering the nature of the Gaussian filter used in the image adapter, the results suggest that images from DTD and FGVCAircraft are situated at two distinct extremes. One necessitates preserving high-frequency details and edge features, while the other calls for more extensive smoothing to alleviate noise. As the sigma decreases, APLe enhances performance in accuracy for both base and novel classes in the DTD dataset while the increase in sigma improves the generalization capability in FGVCAircraft. In section 4.5, the further characteristics and advantages of APLe are explored in detail to adopt V-L models.
4.3 Cross-Dataset Generalization
We evaluate the generalization performance between datasets from ImageNet to other 10 datasets by training 1000 classes. Table 3 demonstrates the ability among CoOp, Co-CoOp and MaPLe. APLe exhibits a modest result on the source dataset, yet ultimately demonstrated a commonly stronger competitive performance of 66.23% with 6/10 datasets surpassing Co-CoOp and 5/10 datasets surpassing MaPLe. APLe achieves equivalent performance in other datasets and outperforms 3 top records. This indicates that the APLe achieves comparable generalization performance with other methods.
Source Target
ImageNet -V2 -Sketch -A -R Average
CLIP 66.73 60.83 46.15 47.77 73.96 57.18
CoOp 71.51 64.20 47.99 49.71 75.21 59.28
Co-CoOp 71.02 64.07 48.75 50.63 76.18 59.91
MaPLe 70.72 64.07 49.15 50.90 76.98 60.28
APLe 70.55 64.15 49.24 50.59 77.38 60.34
Table 4: Comparison in domain generalization.
4.4 Domain Generlization
By training the model on the ImageNet and evaluating the modal in the out-of-domain datasets, Our approach excels in the generalization performance on out-of-distribution datasets among CoOp, Co-CoOp and MaPLe. The average results of APLe are among the top as indicated in Table 4 argues that deploying the token-wise knowledge training with adaptation for multi-modal prompt supports APLe to attain a competitive generalization capability among the methods.
4.5 Vital Experiments and In-Depth Analysis
In the aforementioned foundational experiments, APLe achieves comparable performance with the current state-of-the-art model. Another significance of APLe is that it promotes synergy for collaboration between language text and vision prompts learning to address the challenge in prompt learning due to its settings. In the following vital experiments and in-depth analysis, APLe stands out with an excellent robust and comprehensive generalization capability to adopt the V-L model compared to the baselines in downstream tasks.
Both Language Prompt Only Vision Prompt Only
Base Novel HM Base Novel HM Base Novel HM
3 82.33 74.20 78.06 81.95 73.00 (77.22) 82.10 74.63 78.18
4 81.60 72.21 76.62 81.24 71.71 (76.18) 82.04 74.70 78.20
5 82.27 73.45 77.61 81.72 73.01 77.12 80.44 72.37 (76.19)
6 82.05 71.86 76.62 81.16 70.10 (75.23) 80.40 71.74 75.83
8 81.94 71.03 76.10 80.75 70.21 (75.11) 81.30 71.67 76.18
10 82.62 71.76 76.81 79.53 70.71 (74.86) 81.83 71.74 76.45
AVG 82.14 72.42 76.97 81.06 71.46 (75.96) 81.35 72.81 76.84
Table 5: Prompt Length Analysis Separated in vision and language.
Prompt Length and Overfitting: We initially explored the performance of vision and language as well as both prompts across different prompt lengths shown in Table. 5 with a simplified structure demonstrated As the prompt length increases for both prompts or solely for the vision prompt, the generalization performances experience minor fluctuations along with a modest decrease. Meanwhile, the increase in text prompt length leads to a noticeable decrease in accuracy. In other words, the increase in prompt length in the language prompt has a greater impact compared to that of the vision prompt while these effects are often characterized by a decrease in accuracy. The impact of decreasing performance is commonly described as overfitting in existing papers multi-modal (17) due to prompt conditions such as length. The results suggest that prompt learning behavior varies across the two different modalities including generalization capability and the risk of knowledge conflicts. Owing to the designed framework, APLe effectively improves the generalization performance and mitigates the adverse influences of challenges from the divergence of modalities token and the overfitting introduced by prompt length grows in training while both challenges are well discussed in existing researches.
We evaluate the generalization capability of APLe over different prompt lengths and introduce two evaluations for generalization evaluation. The harmonic mean indicates the generalization performance as foundational experiments while the standard deviation effectively measures the variability in predictive results across 7 different prompt lengths. The evaluations provide an assessment of overfitting for the models. In Table.6, the average performance of APLe provides significant advantages over MaPLe in both accuracy and standard deviation. The average accuracy over 7 different lengths and its standard deviation to the population are 77.07% and 1.37 for APLe, 76.68% and 1.92 for MaPLe. In Fig. 3, the equivalent outcomes repeat when illustrating the generalization performance in two chosen datasets containing numerous textures and small objects and containing the most categories and fine-grained, EuroSAT and StanfordCars, where APLe indicates the more stable curves in both evaluations while the detailed evaluation is demonstrated. The outcomes show that APLe’s framework assists it in further improving generalization capability over the baselines.
Name Evl. Base Novel HM
MaPLe ACC. 81.88 72.10 76.68
Std. 0.78 2.47 1.92
ACC. 81.86 72.82 77.07
APLe Std. 0.64 1.80 1.37
Table 6: Generalization performance comparison between APLe and MaPLe across various prompt lengths in the average of 11 datasets.
Figure 3: Generalization performance comparison (HM) across various prompt lengths in the datasets, EuroSAT and StanfordCars.
Figure 4: noAdapt $vs$ Adapt Comparison.
Effectiveness of Multi-modal Token Adaptation: The designed framework incorporates CLIP zero-shot knowledge to encourage APLe to learn the knowledge comprehensively and independently while alleviating overfitting challenges caused by prompt settings. However, this potentially leads to difficulties in effective collaboration between two tokens. Thus, we introduce token adaptation to promote the synergy between prompt learning and further enhance the model’s generalization performance. In Fig. 4, the results of the analysis in chosen datasets, DTD, EuroSAT, OxfordPets and ImageNet, as well as average performance over 11 datasets, exhibit significant improvement in base classes and overall performance, which suggests the token adaptation supports APLe in improving the generalization capability during prompt learning and provide advanced performance in downstream tasks.
DTD
OxfordPets
$\lambda_{d}$ $\lambda_{g}$ 0.3 0.4 0.5 0.6 0.7 0.8 AVG
0.5 67.19 65.79 63.61 67.95 66.85 68.65 66.67
0.6 65.21 67.38 67.37 68.13 68.54 68.11 67.46
0.7 67.88 66.92 67.59 67.41 68.13 69.19 67.85
0.8 67.98 68.51 68.15 69.36 69.42 66.73 68.36
AVG 67.07 67.15 66.68 68.21 68.24 68.17 67.58
$\lambda_{d}$ $\lambda_{g}$ 0.3 0.4 0.5 0.6 0.7 0.8 AVG
0.5 96.37 95.93 96.48 96.08 95.79 96.24 96.15
0.6 96.46 96.49 96.41 96.58 96.40 96.25 96.43
0.7 96.20 96.32 96.54 96.27 96.39 96.24 96.33
0.8 96.01 96.40 96.38 96.38 96.26 96.65 96.35
AVG 96.26 96.28 96.45 96.33 96.21 96.34 96.31
Table 7: Effectiveness analysis on the hyper-parameter $\lambda$ for both modalities with two chosen datasets, DTD and OxfordPets.
OxfordPets
Effectiveness of hyper-parameter $\lambda$: We further analyze the effect of the hyper-parameter $\lambda$ ($d$ and $g$) used in cooperating zero-shot knowledge for corresponding modality, language and vision respectively, token training as described in Eq. 1819. Results are shown in Table. 7. As pointed out previously, various feature representation complexities exist between the vision and language modalities. The distinctions are demonstrated explicitly in generalization performance and the risk of knowledge conflicts in extended experiments in prompt length. For the results shown in Table. 7, the accuracy in DTD, another dataset containing textures than EuroSAT, improves with increasing values of $\lambda_{d}$ while the result experiences fluctuations when $\lambda_{g}$ also rise. Alternatively, when $\lambda_{d}$ is in the vicinity of 0.8, accuracy achieves higher values. For OxfordPets with the fewest categories and fine-grained, the result indicates a similar pattern to achieve higher value, while the $\lambda_{d}$ is around 0.6. The result suggests that the prompt in the language modality shows a more stable generalization capability for knowledge learning and knowledge conflict resolution than the prompt in the vision modality.
5 Conclusion
The fundamental concern of prompt learning lies in the sensitivity to the setting, such as prompt lengths, while the extended prompts carry the risk of model overfitting. Likewise, the divergences between modalities contribute to the challenge of learning, which also further heightens the risk of overfitting. To this end, APLe introduced the sequential token-wise knowledge training and adaptation framework to address the challenges of multi-modal prompt learning. The proposed approach achieves comparable performance compared to the current state-of-the-art methods in three fundamental tasks. Particularly in datasets with domain shifts, indicates a significant gain in average accuracy. Of greater importance, the results in the analysis of the prompt length and overfitting exhibit APLe dominate the generalization performance across different prompt lengths in downstream tasks.